OpenStudy (anonymous):

solve for the variable : 5-x=16

jimthompson5910 (jim_thompson5910):

We want to isolate x, so we want to get x to one side and get everything else to the other side

OpenStudy (anonymous):

so we add 5 to both sides ?

jimthompson5910 (jim_thompson5910):

5-x is the same as -x+5

jimthompson5910 (jim_thompson5910):

so 5-x = 16 is the same as -x+5 = 16

OpenStudy (anonymous):

so what exactly do i do next

jimthompson5910 (jim_thompson5910):

you want to move that 5 over

jimthompson5910 (jim_thompson5910):

we're adding 5, so we want to undo this and do what?

OpenStudy (anonymous):

not sure

jimthompson5910 (jim_thompson5910):

what's the opposite of addition?

OpenStudy (anonymous):

subtract

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so you subtract 5 from both sides

OpenStudy (anonymous):

which is 11

jimthompson5910 (jim_thompson5910):

good, so -x = 11

jimthompson5910 (jim_thompson5910):

we want x and NOT -x, so we need to get rid of that negative somehow

OpenStudy (anonymous):

how ?

jimthompson5910 (jim_thompson5910):

we can write -x as -1*x

jimthompson5910 (jim_thompson5910):

so -x = 11 is the same as -1*x = 11

jimthompson5910 (jim_thompson5910):

now divide both sides by -1 to completely isolate x -1*x = 11 x = 11/(-1) x = -11

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

glad I could help

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

you're welcome